The area of a rectangle is 48 square centimeters. If the width is 1/3 of the lenght, find the dimensions of the rectangle.
Answer by checkley77(12844)

xy=48
x=y/3
y/3*y=48
y^2/3=48
y^2=3*48
y^2=144
y=sqrt144
y=12 ans.
x=12/3
x=4 ans.
Proof:
4*12=48
48=48